### v1

@@ 3 Discussion
UniCardio introduces a unified framework for multi-modal cardiovascular signal generation. Departing from traditional task-specific methods in cardiovascular signal processing, UniCardio develops a specialized diffusion transformer to capture the multi-modal conditional distributions of cardiovascular signals, including PPG, ECG, and BP, enabling versatile signal restoration and modality translation. By integrating modality-specific encoders and decoders, task-specific attention masks, and a continual learning paradigm, UniCardio achieves superior performance across a broad spectrum of generation tasks with diverse input-output configurations. These capabilities offer considerable benefits in improving monitoring applicability and diagnostic performance for healthcare applications. The model’s generative nature further allows human experts to validate the generated signals, analyze their evolution, and ensure alignment with clinical expectations. UniCardio also demonstrates remarkable efficiency in parameter costs and inference times, making it promising for deployment in wearable devices.
@@
Cardiovascular diseases often develop silently over extended periods, with sudden acute events posing significant health risks. Continuous monitoring in daily life is essential to detect early warning signs and facilitate timely medical intervention. UniCardio enables accurate data collection via adaptive signal restoration, especially for wearable signals that are susceptible to noise and interruptions. Additionally, certain cardiovascular signals remain inaccessible to wearable sensors, which can be synthesized by UniCardio to provide more comprehensive assessments. For patients with severe health conditions, prolonged clinical monitoring that involves non-wearable and invasive procedures can cause considerable discomfort. In such cases, modality translation offers an effective alternative for real-time alerts that prompt necessary clinical examinations. Additionally, integrating generated data with AI-driven diagnostic models further enhances automated diagnoses and provides clinicians with valuable insights, supporting proactive healthcare management.
UniCardio is also expected to facilitate research in fields such as psychological and cognitive sciences, where physiological signals are pivotal for analyzing stress karthikeyan2011review ; de2011stress , cognitive workload momeni2019real ; ranchet2017cognitive , emotion recognition shu2018review ; ayata2020emotion , etc. In such non-critical care scenarios, where wearable devices serve as the primary means of monitoring, the ability to generate high-fidelity non-wearable signals or even invasive signals from wearable inputs is expected to provide more comprehensive assessments. Compared to the complex ICU data targeted in this work, generating physiological signals for non-critical care scenarios is technically less challenging. As such, UniCardio represents a promising solution for harnessing underutilized physiological signals and expanding their applicability to broader scientific domains.
@@
We emphasize that UniCardio’s full potential can be further unleashed by collecting more pre-training data covering a variety of signal modalities with both normal and abnormal conditions. Due to the limited availability of public datasets, our demonstration uses a modest amount of ICU patient data of three modalities, yet achieves compelling results. Future improvements in model performance are expected by increasing the amount of pre-training data for currently involved signal modalities, incorporating additional relevant signal modalities, or introducing supervised information for abnormal conditions. Notably, UniCardio’s model architecture and training paradigm inherently support its continual refinement.
Looking ahead, UniCardio holds the promise of establishing a foundational framework for AI-assisted healthcare, enabling real-time generation of high-fidelity signals for diverse applications, continually integrating knowledge from heterogeneous datasets, and personalizing solutions to individual needs. We expect subsequent work to further refine these capabilities and broaden the scope of multi-modal generation, driving translational breakthroughs in personalized healthcare, next-generation diagnostics, and population-level health monitoring.
@@ 4.1 Problem Formulation
Signal Restoration involves two typical tasks. One is to remove the undesired information from observed recordings, such as denoising the background noise that affects the signal quality index (SQI) elgendi2016optimal . The other is synthesizing the desired information, such as the imputation of the observed intermittent signals xu2022pulseimpute ; bansal2021missing . Both tasks involve transforming the low-quality data $\tilde{\bm{x}}$ into its high-quality counterpart, modeled by the conditional distribution $p(\bm{x}|\tilde{\bm{x}})$. When multi-modal condition information is available for signal restoration, the conditional distribution changes to $p(\bm{x}|\tilde{\bm{x}},\bm{y})$ or $p(\bm{x}|\tilde{\bm{x}},\bm{y},\bm{z})$.
@@
In summary, the restoration and translation of multi-modal cardiovascular signals can be formulated as an assembly of conditional distributions at both condition and modality levels. Without loss of generality, we first assume the signal of target modality as $\bm{x}$ and the available cardiovascular signals are sampled from $q(\bm{x},\bm{y},\bm{z})$. At the condition level, we then aim to capture the modality-specific conditional distributions $p(\bm{x}|\bm{c_{x}})$, where $\bm{c_{x}}$ includes $12^{k}-$ 1 possibilities: (1) signal restoration $p(\bm{x}|\tilde{\bm{x}})$; (2) cross-modality translation $p(\bm{x}|\bm{y})$ and $p(\bm{x}|\bm{z})$; (3) multi-modal signal restoration $p(\bm{x}|\tilde{\bm{x}},\bm{y})$, $p(\bm{x}|\tilde{\bm{x}},\bm{z})$, and $p(\bm{x}|\tilde{\bm{x}},\bm{y},\bm{z})$; and (4) multi-modal signal translation $p(\bm{x}|\bm{y},\bm{z})$. At the modality level, each signal of target modality (e.g., $\bm{x}$, $\bm{y}$, or $\bm{z}$) corresponds to the $12^{k}-$ 1 possibilities of conditional distributions, expanding the total number of tasks to $k\times(2^{k}-1)$ with $p(\bm{x}|\bm{c_{x}})$, $p(\bm{y}|\bm{c_{y}})$, and $p(\bm{z}|\bm{c_{z}})$.
@@ 4.2 Generative Framework
Moreover, UniCardio scales to more signal modalities, e.g., at least three modalities (PPG, ECG, and BP) versus two modalities (text and image). Providing modality-specific time steps to diffusion models combinatorially increases the complexity of generation tasks, resulting in limited training efficiency and overall performance. To address this, we propose a novel design that simplifies the several modality-specific time steps bao2023one to a unified time step for all modalities, e.g., from $p_{\bm{\theta}}(\bm{y}_{t-1}|\bm{x}_{t^{x}},t^{x},\bm{y}_{t^{y}},t^{y},\bm{z}_{t^{z}},t^{z})$ to $p_{\bm{\theta}}(\bm{y}_{t-1}|\bm{x}_{t},\bm{y}_{t},\bm{z}_{t},t)$. This unified time step enables a seamless integration of multi-modal cardiovascular signal restoration and translation, as shown in Eq. (1) and Eq. (2). Further details are provided in Supplementary Sec. B.
@@ 4.3 Model Architecture
The task-specific attention mask $M$ controls the information flow with predefined values. Each modality $i\in[k]$ maps to the token range $[(i-1)L:iL]$ of $\bm{h}$ (Supplementary Table 4). For translation tasks from modality $i$ to modality $j$, $M$ permits intra-modal interactions within $i$, intra-modal interactions within $j$, and inter-modal interactions from $i$ to $j$. The elements of $M$ corresponding to these interactions are set to zero, while other elements are assigned large negative values to block irrelevant interactions during the softmax operation. For denoising or imputation tasks, where AM acts as the target modality ($j=k$), $M$ similarly ensures information flow of intra-modal interactions within $i$, intra-modal interactions within $k$, and inter-modal interactions from $i$ to $k$, while blocking other irrelevant interactions. This unified masking mechanism enables consistent handling of generative tasks involving different modalities and input-output configurations.
The output $\bm{h}^{{}^{\prime}}$ from the MSA layer is refined by a fully connected (FC) layer $F_{1}(\cdot)$ and combined with the time embedding $\bm{h}_{t}$, updating $\bm{h}^{{}^{\prime}}\leftarrow F_{1}(\bm{h}^{{}^{\prime}})+\bm{h}_{t}$. This operation expands the feature dimension, resulting in a feature vector of shape $(B,kL,2C)$. We then implement a gated activation unit van2016conditional ; van2016wavenet to capture the complex conditional distributions of signal modalities given the updated feature vector $\bm{h}^{{}^{\prime}}$. Specifically, we split $\bm{h}^{{}^{\prime}}$ along the feature dimension into two parts, i.e., $\bm{h}^{{}^{\prime}}=[\bm{h}^{{}^{\prime}}_{1};\bm{h}^{{}^{\prime}}_{2}]$, each of shape $(B,kL,C)$. These two parts are processed separately through $\tanh$ and $\sigma$ activations, respectively, and combined element-wise as $\bm{h}^{\prime}\leftarrow\tanh(\bm{h}^{\prime}_{1})\odot\sigma(\bm{h}^{\prime}_{2})$. This operation produces a transformed feature vector $\bm{h}^{{}^{\prime}}$ with a reduced dimension $(B,kL,C)$.
We further implement residual and skip connections kongdiffwave ; tashiro2021csdi across customized transformer modules to facilitate convergence of the entire model. Specifically, the transformed feature vector $\bm{h}^{{}^{\prime}}$ is projected back to an expanded dimension of $(B,kL,2C)$ though another FC layer $F_{2}(\cdot)$ and split again into two parts, $\bm{h}^{{}^{\prime}}=[\bm{h}^{{}^{\prime}}_{1};\bm{h}^{{}^{\prime}}_{2}]$, each of shape $(B,kL,C)$. The first part $\bm{h}^{{}^{\prime}}_{1}$ is added to the input of the current transformer module, serving as the input to the next transformer module. The second part $\bm{h}^{{}^{\prime}}_{2}$ is accumulated across all transformer modules and passed through an additional FC layer $F_{3}(\cdot)$. This produces the final feature vector $\bm{h}^{{}^{\prime}}_{s}=F_{3}(\sum\bm{h}^{{}^{\prime}}_{2})$, retaining the shape $(B,kL,C)$.
@@
In particular, the first $k-1$ decoders $D_{i},i\in[k-1]$ are optimized from network training, whereas the last decoder $D_{k}$ of AM inherits weights from one of the $k-1$ decoders depending on the generation tasks. For example, when performing ECG imputation with $k=4$ representing PPG, ECG, BP and AM, the decoder $D_{k}$ reuses the weights of $D_{2}$ to generate ECG signals. This strategy ensures consistency across tasks and reduces the need for additional training of the final decoder.
@@ 4.4 Experimental Setup
Dataset. UniCardio is pre-trained with the Cuffless BP dataset kachuee2016cuffless , which contains 339 hours of trimodal recordings (PPG, ECG, and BP) collected from ICU patients. PPG and ECG signals are processed using a 0.5 Hz high-pass Butterworth filter (order = 5) and a 50 Hz pulse filter, followed by z-score normalization makowski2021neurokit2 . BP signals are left unfiltered and unnormalized to preserve their absolute magnitude. To ensure data quality, we address recording errors such as sensor displacements and device disconnections by filtering out noisy signals based on sample entropy. This metric quantifies the regularity and complexity of time series by measuring the likelihood that a given pattern remains consistent in subsequent points. We set thresholds of 0.2, 0.3, and 0.2 for PPG, ECG, and BP signals, respectively, to exclude highly irregular or complex signals. ECG signals are more prone to corruption due to their high-frequency components and therefore require additional pre-processing. We first detect and rectify inverted ECG signals caused by sensor displacements, and then apply SQI-based selection makowski2021neurokit2 for further refinements. Finally, PPG and ECG signals are min-max normalized, while BP signals are z-scored using the mean and standard deviation calculated from the training set. The processed signals are segmented into 4-second pairs and split into training, validation, and test sets in an 80%-10%-10% ratio.
@@
The task-specific attention masks control the condition modalities for each generation task throughout the training process. The model is optimized using an SGD optimizer over the total $e$ epochs. The learning rate starts as $1\times 10^{-3}$ at the beginning of the first $e/k$ epochs to provide a strong initialization for one-condition tasks. It is then reduced to $1\times 10^{-4}$ at epoch $0.7\times e/k$ as the condition modalities increases. In the final $e/k$ epochs, the learning rate is further decreased to $1\times 10^{-5}$ to enable precise fine-tuning and ensure balanced performance across all tasks. We empirically identify that $e=800$ already achieves superior performance. Further details are provided in Supplementary Sec. C, and the pseudo codes are provided in Supplementary Sec. E.
Downstream Application. After the pre-training stage, we consider a variety of downstream applications on different datasets, including the PTBXL dataset wagner2020ptb for detection of abnormal health conditions, the MIMIC PERform AF dataset charlton2022detecting for detection of AF, the WESAD dataset schmidt2018introducing for HR estimation, and the unseen data from the Cuffless BP dataset kachuee2016cuffless for BP estimation. These datasets are processed with a similar pipeline as the pre-training dataset. The detection of AF, ST change, and hypertrophy is performed by training respective classification models based on 1D VGG-16 architectures liu2022multiclass . HR estimation is conducted by detecting heartbeat peaks with common algorithms nabian2018open . BP estimation is achieved by training a regression model based on a CNN-LSTM architecture jeong2021combined .
Evaluation Metric. We evaluate the quality of generated signals with three common metrics, including the root mean squared error (RMSE), mean absolute error (MAE), and Kolmogorov-Smirnov test (KS-Test). Additionally, we assess the ability to detect cardiovascular abnormalities with accuracy, specificity, and sensitivity, which reflect different proportions of true positives (TP), true negatives (TN), false positives (FP), and false negatives (FN).
@@
Our code will be released at https://github.com/thu-ml/UniCardio.

### v2

@@ 3 Discussion
UniCardio introduces a unified framework for multi-modal cardiovascular signal generation. Departing from traditional task-specific methods in cardiovascular signal processing, UniCardio develops a specialized diffusion transformer to model multi-modal conditional distributions of cardiovascular signals, including PPG, ECG, and BP, enabling versatile signal restoration and modality translation. By integrating modality-specific encoders and decoders, task-specific attention masks, and a continual learning paradigm, UniCardio achieves superior performance across a broad spectrum of generation tasks with diverse input-output configurations. These capabilities significantly enhance monitoring applicability and diagnostic performance for healthcare applications. The model’s generative nature further allows human experts to validate the generated signals, analyze their evolution, and ensure alignment with clinical expectations. UniCardio also demonstrates remarkable efficiency in parameter costs and inference times, making it promising for deployment in wearable devices.
@@
Cardiovascular diseases often develop silently over extended periods, with acute events posing significant health risks. Continuous monitoring in daily life is essential to detect early warning signs and facilitate timely medical intervention. UniCardio enables accurate data collection via adaptive signal restoration, especially for wearable signals that are susceptible to noise and interruptions. Additionally, certain cardiovascular signals remain inaccessible to wearable sensors, which can be synthesized by UniCardio to provide more comprehensive assessments. For patients with severe health conditions, prolonged clinical monitoring that involves non-wearable and invasive procedures can cause considerable discomfort. In such cases, modality translation offers an effective alternative for real-time alerts that prompt necessary clinical examinations. Integrating generated data with AI-driven diagnostic models further enhances automated diagnoses and offers clinicians meaningful insights, supporting proactive healthcare management.
Beyond healthcare, UniCardio may also benefit research in psychological and cognitive sciences, where physiological signals are widely used to assess stress karthikeyan2011review ; de2011stress , cognitive workload momeni2019real ; ranchet2017cognitive , and emotion recognition shu2018review ; ayata2020emotion . In particular, pulse wave amplitude from PPG xuan2020assessing ; pavlov2023task , HR and HR variability from ECG al2018objective ; mandrick2016neural , as well as SBP and DBP richter2008task ; el2016rate are common indicators of mental effort during cognitive tasks. In such non-critical care scenarios, wearable devices are typically the primary monitoring tools but often yield recordings with greater interruptions than well-controlled ICU environments. UniCardio is expected to address the recording challenges through its signal restoration capabilities. Conversely, physiological signals collected from generally healthy subjects in psychological and cognitive studies tend to exhibit higher regularity and fewer pathological features compared to ICU patient data, potentially simplifying generation tasks under certain conditions. Although empirical validation remains necessary, UniCardio’s demonstrated performance holds promise for improving data quality and analytical depth in these scientific domains.
@@
We acknowledge several limitations that warrant consideration. First, UniCardio is pre-trained on a modest amount of ICU patient data comprising three signal modalities due to the limited availability of large-scale, multi-modal public datasets. While the model achieves compelling results across a range of tasks and benchmarks, future improvements may benefit from expanded pre-training data, additional signal modalities, and annotated health conditions. Notably, UniCardio’s model architecture and training paradigm inherently support continual refinement as new data and modalities become available. Second, while UniCardio shows promising results in handling noisy and incomplete signals, including those resembling wearable-device recordings, its current evaluations are based on benchmark datasets collected in controlled clinical settings. As such, caution is warranted when interpreting its performance for real-world wearable applications, where motion artifacts, sensor variability, and environmental interference may introduce additional challenges. These scenarios require empirical validation and potentially domain-specific adaptation.
Looking ahead, UniCardio offers a versatile and scalable foundation for multi-modal physiological signal generation, with demonstrated strengths in both signal restoration and modality translation across diverse configurations. Its model architecture and training paradigm are inherently designed to support continual integration of new datasets, signal modalities, and task demands. Future research is expected to extend UniCardio’s capabilities toward broader domains of AI-assisted healthcare, including personalized monitoring, automated diagnostics, and integrated physiological data analysis.
@@ 4.1 Problem Formulation
[…]
@@
[…]
@@ 4.2 Generative Framework
Moreover, UniCardio scales to more signal modalities, e.g., at least three modalities (PPG, ECG, and BP) versus two modalities (text and image). Providing modality-specific time steps to diffusion models combinatorially increases the complexity of generation tasks, resulting in limited training efficiency and overall performance. To address this, we propose a specialized design that simplifies the several modality-specific time steps bao2023one to a unified time step for all modalities, e.g., from $p_{\bm{\theta}}(\bm{y}_{t-1}|\bm{x}_{t^{x}},t^{x},\bm{y}_{t^{y}},t^{y},\bm{z}_{t^{z}},t^{z})$ to $p_{\bm{\theta}}(\bm{y}_{t-1}|\bm{x}_{t},\bm{y}_{t},\bm{z}_{t},t)$. This unified time step enables a seamless integration of multi-modal cardiovascular signal restoration and translation, as shown in Eq. (1) and Eq. (2). Further details are provided in Supplementary Sec. B.
@@ 4.3 Model Architecture
The task-specific attention mask $M$ controls the information flow with predefined values. Each modality $i\in[k]$ maps to the token range [(i1)L:iL][(i-1)L:iL] of $\bm{h}$ (Supplementary Table 5). For translation tasks from modality $i$ to modality $j$, $M$ permits intra-modal interactions within $i$, intra-modal interactions within $j$, and inter-modal interactions from $i$ to $j$. The elements of $M$ corresponding to these interactions are set to zero, while other elements are assigned large negative values to block irrelevant interactions during the softmax operation. For denoising or imputation tasks, where AM acts as the target modality ($j=k$), $M$ similarly ensures information flow of intra-modal interactions within $i$, intra-modal interactions within $k$, and inter-modal interactions from $i$ to $k$, while blocking other irrelevant interactions. This unified masking mechanism enables consistent handling of generative tasks involving different modalities and input-output configurations.
The output $\bm{h}^{{}^{\prime}}$ from the MSA layer is refined by a fully connected (FC) layer $F_{1}(\cdot)$ and combined with the time embedding $\bm{h}_{t}$, updating $\bm{h}^{{}^{\prime}}\leftarrow F_{1}(\bm{h}^{{}^{\prime}})+\bm{h}_{t}$. This operation expands the feature dimension, resulting in a feature vector of shape $(B,kL,2C)$. We then implement a gated activation unit van2016conditional ; van2016wavenet to capture the complex conditional distributions of signal modalities given the updated feature vector $\bm{h}^{{}^{\prime}}$. Specifically, we split $\bm{h}^{{}^{\prime}}$ along the feature dimension into two parts, i.e., $\bm{h}^{{}^{\prime}}=[\bm{h}^{{}^{\prime}}_{1}\semicolon\;\bm{h}^{{}^{\prime}}_{2}]$, each of shape $(B,kL,C)$. These two parts are processed separately through $\tanh$ and $\sigma$ activations, respectively, and combined element-wise as $\bm{h}^{\prime}\leftarrow\tanh(\bm{h}^{\prime}_{1})\odot\sigma(\bm{h}^{\prime}_{2})$. This operation produces a transformed feature vector $\bm{h}^{{}^{\prime}}$ with a reduced dimension $(B,kL,C)$.
We further implement residual and skip connections kongdiffwave ; tashiro2021csdi across customized transformer modules to facilitate convergence of the entire model. Specifically, the transformed feature vector $\bm{h}^{{}^{\prime}}$ is projected back to an expanded dimension of $(B,kL,2C)$ though another FC layer $F_{2}(\cdot)$ and split again into two parts, $\bm{h}^{{}^{\prime}}=[\bm{h}^{{}^{\prime}}_{1}\semicolon\;\bm{h}^{{}^{\prime}}_{2}]$, each of shape $(B,kL,C)$. The first part $\bm{h}^{{}^{\prime}}_{1}$ is added to the input of the current transformer module, serving as the input to the next transformer module. The second part $\bm{h}^{{}^{\prime}}_{2}$ is accumulated across all transformer modules and passed through an additional FC layer $F_{3}(\cdot)$. This produces the final feature vector $\bm{h}^{{}^{\prime}}_{s}=F_{3}(\sum\bm{h}^{{}^{\prime}}_{2})$, retaining the shape $(B,kL,C)$.
@@
In particular, the first k1k-1 decoders $D_{i},i\in[k-1]$ are optimized from network training, whereas the last decoder $D_{k}$ of AM inherits weights from one of the k1k-1 decoders depending on the generation tasks. For example, when performing ECG imputation with $k=4$ representing PPG, ECG, BP and AM, the decoder $D_{k}$ reuses the weights of $D_{2}$ to generate ECG signals. This strategy ensures consistency across tasks and reduces the need for additional training of the final decoder.
@@ 4.4 Experimental Setup
Dataset. UniCardio is pre-trained with the Cuffless BP dataset kachuee2016cuffless , which contains 339 hours of trimodal recordings (PPG, ECG, and BP) collected from ICU patients. PPG and ECG signals are processed using a 0.5 Hz high-pass Butterworth filter (order = 5) and a 50 Hz pulse filter, followed by z-score normalization makowski2021neurokit2 . BP signals are left unfiltered and unnormalized to preserve their absolute magnitude. To ensure data quality, we address recording errors such as sensor displacements and device disconnections by filtering out such signals based on sample entropy richman2000physiological , which quantifies the regularity and complexity of time series by estimating the likelihood that similar patterns remain consistent over time. We empirically determined entropy thresholds by analyzing the distribution of values across the training set and qualitatively inspecting corresponding signal segments. Specifically, we found that segments with entropy values above 0.2 for PPG and BP, and above 0.3 for ECG, typically exhibit excessive irregularity or noise and lack the expected periodic or quasi-periodic structure. These thresholds effectively remove physiologically implausible segments while preserving sufficient high-quality data for training and evaluation. ECG signals are more prone to corruption due to their high-frequency components and therefore require additional pre-processing. We first detect and rectify inverted ECG signals caused by sensor displacements, and then apply SQI-based selection makowski2021neurokit2 for further refinements. Finally, PPG and ECG signals are min-max normalized, while BP signals are z-scored using the mean and standard deviation calculated from the training set. The processed signals are segmented into 4-second pairs and randomly split into training, validation, and test sets in an 80%-10%-10% ratio.
@@
The task-specific attention masks control the condition modalities for each generation task throughout the training process. The model is optimized using an SGD optimizer over the total $e$ epochs. The learning rate starts as $1\times 10^{-3}$ at the beginning of the first $e/k$ epochs to provide a strong initialization for one-condition tasks. It is then reduced to $1\times 10^{-4}$ at epoch $0.7\times e/k$ as the condition modalities increases. In the final $e/k$ epochs, the learning rate is further decreased to $1\times 10^{-5}$ to enable precise fine-tuning and ensure balanced performance across all tasks. We empirically identify that $e=800$ already achieves superior performance. Further details are provided in Supplementary Sec. C, and the pseudo codes are provided in Supplementary Sec. E. The entire model is pre-trained with 8-card RTX 4090 GPUs for around 7 days. The carbon footprint is around 225kg CO2.
Downstream Application. After the pre-training stage, we consider a variety of downstream applications on different datasets, including the PTBXL dataset wagner2020ptb for detection of abnormal health conditions, the MIMIC PERform AF dataset charlton2022detecting for detection of AF, the WESAD dataset schmidt2018introducing for HR estimation, and the MIMIC dataset moody1996database for SBP/DBP estimation. These datasets are processed with a similar pipeline as the pre-training dataset. The detection of AF, ST change, and hypertrophy is performed by training respective classification models based on 1D VGG-16 architectures liu2022multiclass . HR estimation is conducted by detecting heartbeat peaks with common algorithms nabian2018open . SBP/DBP estimation is conducted by training a regression model based on a CNN-LSTM architecture jeong2021combined .
Evaluation Metric. We evaluate the quality of generated signals with three common metrics, including the root mean squared error (RMSE), mean absolute error (MAE), and Kolmogorov-Smirnov test (KS-Test). To account for small temporal misalignments that may disproportionately affect pointwise metrics, we further report the minimum RMSE (Min-RMSE) and minimum MAE (Min-MAE) based on bounded temporal shifts. These metrics quantify the best achievable reconstruction error after applying small temporal shifts to generated signals, providing a more robust assessment of morphological similarity. Moreover, we assess the ability to detect cardiovascular abnormalities with accuracy, specificity, and sensitivity, which reflect different proportions of true positives (TP), true negatives (TN), false positives (FP), and false negatives (FN).
@@
  • Min-RMSE quantifies the lowest achievable RMSE between generated and ground-truth signals after applying small temporal shifts to account for misalignments:
    $\mathrm{Min\text{-}RMSE}=\min_{\tau\in\mathcal{T}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\hat{\bm{x}}^{(\tau)}_{i}-\bm{x}_{i}\right)^{2}},$ (9)
    where $\mathcal{T}$ denotes a finite set of allowable temporal shifts (within one second in our case), and $\hat{\bm{x}}^{(\tau)}_{i}$ correspond to $\hat{\bm{x}}_{i}$ shifted by $\tau$ signal points.
  • Min-MAE similarly reports the minimum achievable MAE after applying small temporal shifts:
    $\mathrm{Min\text{-}MAE}=\min_{\tau\in\mathcal{T}}\frac{1}{N}\sum_{i=1}^{N}|\hat{\bm{x}}^{(\tau)}_{i}-\bm{x}_{i}|.$ (10)
@@
The implementation code is included in Supplementary Materials for examination, and will be released at https://github.com/thu-ml/UniCardio upon acceptance.